OpenStudy (anonymous):

Find all the solutions for theta, given 0< or equal to theta < 2pi, for the function Sin 2(theta)-Cos(theta) = 0

OpenStudy (anonymous):

cos(theta)=0 or sin(theta)=(1/2)

OpenStudy (anonymous):

start with \(\sin(2x)=2\sin(x)\cos(x)\)

OpenStudy (anonymous):

theta=pi/2, 3pi/2, pi/6, 5pi/6

OpenStudy (anonymous):

then \[2\sin(\theta)\cos(\theta)-\cos(\theta)=0\] \[\cos(\theta)(2\sin(\theta)-1)=0\] \[\cos(\theta)=0\iff \theta =\frac{\pi}{2},\text{ or } \theta =\frac{3\pi}{2}\]

OpenStudy (anonymous):

or \[2\sin(\theta)-1=0\] \[\sin(\theta)=\frac{1}{2}\] what @josiahh sais

OpenStudy (anonymous):

Thanks guys